# Large Language Models Can Learn Temporal Reasoning

### v2

@@ Abstract
While large language models (LLMs) have demonstrated remarkable reasoning capabilities, they are not without their flaws and inaccuracies. Recent studies have introduced various methods to mitigate these limitations. Temporal reasoning (TR), in particular, presents a significant challenge for LLMs due to its reliance on diverse temporal expressions and intricate contextual details. In this paper, we propose TG-LLM, a new framework towards language-based TR. To be specific, we first teach LLM to translate the context into a temporal graph (TG). A synthetic dataset, which is fully controllable and requires minimal supervision, is constructed for fine-tuning on this graph translation task. We confirm in experiments that the capability of TG extraction learned on our dataset can be transferred to other TR tasks and benchmarks. On top of that, we guide LLM to perform symbolic reasoning over the TG via Chain of Thoughts (CoTs) bootstrapping and special data augmentation strategies. We observe that CoTs with symbolic reasoning bring more consistent and reliable results than those using free-form text.
@@ 1 Introduction
As one of the fundamental abilities, temporal reasoning (TR) plays an important role in human perception. It is not just about understanding basic concepts such as order or duration; it extends to more intricate aspects, e.g., making sense of complex narratives or discovering causal relations. Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI et al., 2023; Touvron et al., 2023b) have shown remarkable performance in complex reasoning tasks  (Huang and Chang, 2023). However, they still can not perform TR sufficiently well  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023). TR is more challenging since it requires a combination of various skills including mathematical and logical reasoning as well as commonsense knowledge  (Zhou et al., 2019; Qin et al., 2021).
@@
Figure 1: Our TG-LLM framework has two steps: text-to-temporal graph translation and neuro-symbolic reasoning.
@@
Recent works mainly adopt the general reasoning methods to investigate and improve the TR capability of LLMs. For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark leading LLMs on different TR tasks with standard prompting, few-shot learning and Chain of Thoughts (CoTs)  (Wei et al., 2023a).  (Li et al., 2023; Yuan et al., 2023) introduce pre-defined Python programs/rule-based templates to perform supervised fine-tuning (SFT). Similarly,  (Wei et al., 2023b) design several special types of prompts for in-context learning (ICL). In addition,  (Tan et al., 2023a, b) adopt some extra strategies, which include specific pre-training, instruction tuning and reinforcement learning.
@@
Despite the effectiveness of such methods, they either ignore or not explicitly involve the intrinsic nature of TR. Human performs complex TR on a timeline of events which are aligned with the entities and relations. These temporal concepts (e.g., ordering, duration, frequency, typical time) are then rigorously defined based on the timeline information. Thus, a natural solution for LLMs to reach human-level performance is to first extract a timeline of events from the text, and then perform deliberate reasoning on it. However, the high-quality extraction of the aligned timeline is a challenging task itself due to the lack of ground truth. To solve this problem, we propose a synthetic dataset, which is fully controllable and requires minimal supervision. We show in experiments that the capability of timeline extraction learned on our dataset can be transferred to other tasks and benchmarks.
@@
Given a reliable timeline, we still need to solve two problems to guide LLMs to perform rigorous reasoning on it: (1) How can one introduce the necessary arithmetic and commonsense knowledge involved in TR? (2) How can one guide LLMs to perform rigorous reasoning? Inspired by retrieval-augmented generation (RAG)  (Lewis et al., 2021), we first extract all the time expressions, and then generate all the necessary information (e.g., the temporal relation and time gap between the timestamps, the relative order of the gaps). As for (2), there are two roadmaps: (i) translating natural language into logical statements, and using external symbolic engine for reasoning  Pan et al. (2023); (ii) directly using LLMs as the reasoning engine  (Zhu et al., 2023). For (i), the difficulty lies in accurate translation  Yang et al. (2023c) and the limited expressive power of formal logic. For (ii), there is no guarantee for the correctness of the generated text especially when training data are insufficient  (Yang et al., 2023b). To solve these problems, we propose a symbolic-integration bootstrapping method to generate reliable intermediate steps for supervised fine-tuning. Furthermore, we design special data augmentation strategies to mitigate data deficiency for neuro-symbolic reasoning tasks.
@@
  • We propose a new paradigm, TG-LLM, for language-based TR. In this framework, we first translate the context into a temporal knowledge graph, and then perform symbolic reasoning on it. Extensive experiments prove that our novel approach results in superior performance compared to the conventional approach.
  • We design two approaches including chain of thoughts bootstrapping and data augmentation to guide LLM to utilize symbolic reasoning, which brings better performance than free-form text.
  • We present a pipeline to create a synthetic dataset for question answering that requires TR. It is fully controllable and requires minimal supervision for text-temporal graph alignment. We show in experiments that fine-tuning on our dataset benefits LLMs on other TR tasks and benchmarks.
@@ 2 Dataset Construction
Relation <Event_B>? _and_B_start_time>
@@
Facts When did the <Event_A> occur? <Event_A_start_time>
@@
Table 2: Reasoning types with corresponding questions and answers in TGQA.
@@
We propose a pipeline for the construction of our synthetic dataset called TGQA. Each sample, as shown in Table 1, is in the form of (story, temporal graph, questions, answers). The temporal graph (TG) is a subgraph extracted from a large temporal knowledge graph, YAGO11k (Dasgupta et al., 2018), with a restriction on the number of events. Each story is translated from the TG by GPT-3.5. Finally, by using some rule-based templates, we obtain reliable question and answer (QA) pairs from the graph. We obtain 500 samples for training/validation and another 100 for test, with about 30 QA pairs in a single sample. Compared with existing datasets  (Chen et al., 2021; Tan et al., 2023a), we have more diverse categories and types of temporal questions as shown in Table 2. Below we provide the details of each step in the dataset construction pipeline.
@@
Step 1: Graph Splitting & Masking  Existing temporal knowledge graph datasets (Leetaru and Schrodt, 2013; Dasgupta et al., 2018; García-Durán et al., 2018) have tens of thousands of events, making it impossible to convert them into text all at once. Given a specific entity, we extract its third-order neighbors and all the events happening between them. Since we want LLMs to do reasoning instead of memorization, it is ensured that no overlapping exists between the facts in training/validation and test samples. In addition, we consider the noticeable corpus leakage problem of LLMs, where prior knowledge of test data has been implicitly obtained from the extensive pre-training process. Thus, a masking strategy, i.e., changing entity names into random ones of the same type, is adopted. We generate a global mapping for entity names with GPT-3.5 (see Table 6). To avoid confusion, we only choose obscure names that do not exist in the original large knowledge graph.
@@
Step 2: Graph-based Open QA Creation  Based on the given TG, we generate a story and multiple QAs. We first ask GPT-3.5 to write a story based on the graph with the requirements to include all the facts. It is observed that GPT-3.5 tends to ignore the end time of some events in the created story. To solve this problem, we split the starting and ending of a persistent event into two transient ones (see Table 1). To obtain a comprehensive benchmark, we consider all types of temporal reasoning, which include sequencing, duration, frequency, simultaneity, temporal relation, comparative analysis and facts extraction. Given these categories, we design multiple diverse question types (see Table 2), and use a rule-based Python script to generate the corresponding Qs and As.
@@
Step 3: Quality Control  Although LLMs are good at story generation, there is a risk of misalignment between the given TG and the generated text. To avoid the interference caused by misalignment, we propose a semi-automatic way for verification. Given a generated story, we ask LLMs the time of each event in the graph. If LLMs cannot give the correct answer, we consider the event as possibly missing in the story. Fully manual inspection is expensive, but by first asking LLM we can narrow down potential errors and only manually inspect the set of unanswered questions by LLM. Note that only the quality of story-TG alignment needs manual inspection, since all the QAs are generated from the TG with specific rules. We show dataset statistics in Appendix A, and all the prompts involved in Appendix C.
@@ 3 TG-LLM
Motivated by human perception, we propose a new paradigm called TG-LLM. We first translate the text into a temporal graph, and then guide the LLM to perform symbolic reasoning on it.
@@ 3.1 Text-to-TG Translation
TG serves as the foundation of the following symbolic reasoning process. Although LLMs can extract fact/events from the given context without fine-tuning, we observe a misalignment between the generated graph and pre-defined QAs, i.e, LLMs making mistakes or using different entity/relation names or time expressions. Thus, we provide a deliberate pipeline for high-quality TG translation.
@@
Ground-truth TG Generation   For datasets such as TGQA, each story is corresponding to a verified TG. Thus, we can directly fine-tune LLMs on this task with the ground truth. However, the biggest challenge for most real-world applications is the lack of ground truth. Instead of using the naive estimation from LLMs, we provide a pipeline for high-quality generation.
(1) Entity & Relation Extraction: To align the entity/relation names in generated TGs and pre-defined QAs, we use a prompt with requirements on including the ones in pre-defined QAs when extracting from the context. To obtain the entity/relation names in pre-defined QAs, we consider two ways: parsing with rules or using LLMs.
@@
(2) Temporal Info Identification: To identify all the time expressions, LLMs are first used to extract them from the context, and then a rule-based Python script is adopted for time normalization and invalid output filtering. For the valid normalized time expressions, we can easily obtain their temporal relations and time gaps, which are introduced as external knowledge in temporal reasoning (see Table 7).
@@
(3) Graph Construction & Verification: We generate the graph using LLMs with ICL. To be specific, we provide in the prompt several examples which include a story and extracted entities, relations and time expressions as input and the corresponding TG as output (see Table 15). For those without explicit expressions, we ask LLMs to provide the temporal order of different events. We further semi-automatically verify the generated TGs by asking LLMs to answer the pre-defined QAs based on the graph. For those graphs where LLMs give the wrong answers, we manually check and correct them.
@@
Figure 2: In chain of thoughts (CoTs) bootstrapping, we accept CoTs leading to correct answers and sample them according to their contrastive learning scores.
@@
Supervised Fine-tuning  We found out in our experiments that providing a chronological list of events helps LLMs provide better temporal reasoning. In addition, as shown in Table 1, to deal with both transient and persistent events, we consider the starting and ending of a persistent event as two transient events. With these steps, we transfer the TG into a timeline with the alignment between entities, relations and times. We perform supervised fine-tuning (SFT) using Llama-2 model  (Touvron et al., 2023a) with Low-Rank Adaptation (LoRA)  (Hu et al., 2021). The input and output of the model are the story and aligned timeline, respectively. In experiments, we observe benefits from the SFT on our dataset to other temporal reasoning tasks.
3.2 Neuro-symbolic Temporal Reasoning
@@
Given the reliable TGs, we train LLMs to perform symbolic reasoning with SFT enhanced by CoTs bootstrapping and data augmentation.
@@ 3.2.1 Bootstrapping Chain of Thoughts
It is observed that SFTs on reliable CoTs bring better performance in reasoning tasks than those on standard Input/Output prompts  (Wang et al., 2023; Ho et al., 2023). However, it is challenging to obtain reliable CoTs from LLMs for symbolic reasoning tasks  (Saparov and He, 2023). Inspired by  (Zhu et al., 2023), we provide several manually verified examples to LLMs as ICL, and ask LLMs to generate CoTs for the new samples.
To be specific, we first classify the pre-defined QAs into several categories $\{Q_{1},Q_{2},\cdots,Q_{M}\}$. For each category $Q_{i}$, we randomly choose $N_{i}$ samples $\{(g_{j},e_{j},q_{j},a_{j})\}^{N_{i}}_{j=1}$ and ask GPT-3.5/GPT-4 to provide symbolic CoTs $\{c_{j}\}$ given the TG $\{g_{j}\}$, external knowledge $\{e_{j}\}$ and question $\{q_{j}\}$. These CoTs will be manually verified and used as few-shot ICL examples. When given new samples $(g_{j^{\prime}},e_{j^{\prime}},q_{j^{\prime}},a_{j^{\prime}})$, we bootstrap CoTs with final answers $(c_{j^{\prime},k},\hat{a}_{j^{\prime},k})$ from GPT-3.5, and only accept the one leading to correct answers, i.e., $\hat{a}_{j^{\prime},k}=a^{*}_{j^{\prime}}$. In addition, inspired by  (Wang et al., 2023), we compare different CoTs using contrastive learning given in Eq. 2. The core idea is to choose CoTs that make the correct answer more plausible than other candidate answers. For the generated CoTs, we use weighted sampling according to the score-based probability given in Eq. 1.
@@
where $P_{sample}(\cdot)$ is the sampling probability for CoTs selection, $score(\cdot)$ is the contrastive learning score for each CoT, $g,e,q,c$ denote the TG, external knowledge, question, and CoT, respectively; $a^{*}$ denotes the correct answer; $a^{\prime}$ denotes the wrong answer from a candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $G(\cdot)$ denotes the contrastive plausibility growth; conditional probability $P(\cdot)$ is calculated in a way similar to perplexity, $t_{l}$ denotes the $l$-th token in the answer $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.
Compared with generating CoTs with templates  (Saparov and He, 2023; Yuan et al., 2023), our method has more diversity and does not require prior knowledge for the reasoning process.
3.2.2 Data Augmentation
Compared with conventional tasks, reasoning suffers more from data insufficiency since more information, such as evidence, arguments, and logics, are involved in intermediate steps  (Huang and Chang, 2023). To address the issue, we propose special strategies for data augmentation.
Essentially, our framework involves two steps: text-to-TG translation and symbolic temporal reasoning. At the test stage, reasoning happens on the noisy estimated graphs, which impacts the overall performance. Specifically, for the reasoning module, we train with ground-truth/verified graphs while in the test time, the estimated graph is used for symbolic reasoning. To improve the robustness of reasoning, we introduce some disturbances on the graphs during training. Note that the type of disturbances should be carefully designed in order to avoid confusing the model. We first investigate two types of disturbances: (i) remove irrelevant edges and (ii) replace edges by changing relation names, given in Eq. 5 and Eq. 6, respectively. An edge is considered irrelevant if it is not involved in both QA and CoT.

### v3

@@ Abstract
While large language models (LLMs) have demonstrated remarkable reasoning capabilities, they are not without their flaws and inaccuracies. Recent studies have introduced various methods to mitigate these limitations. Temporal reasoning (TR), in particular, presents a significant challenge for LLMs due to its reliance on diverse temporal expressions and intricate temporal logic. In this paper, we propose TG-LLM, a novel framework towards language-based TR. Instead of reasoning over the original context, we adopt a latent representation, temporal graph (TG) that facilitates the TR learning. A synthetic dataset (TGQA), which is fully controllable and requires minimal supervision, is constructed for fine-tuning LLMs on this text-to-TG translation task. We confirmed in experiments that the capability of TG translation learned on our dataset can be transferred to other TR tasks and benchmarks. On top of that, we teach LLM to perform deliberate reasoning over the TGs via Chain of Thought (CoT) bootstrapping and graph data augmentation. We observed that those strategies, which maintain a balance between usefulness and diversity, bring more reliable CoTs and final results than the vanilla CoT distillation.  [^1] The code is available at https://github.com/xiongsiheng/TG-LLM.
@@ 1 Introduction
As one of the fundamental abilities, temporal reasoning (TR) plays an important role in human perception. It is not just about understanding basic concepts such as ordering or duration; it extends to more intricate aspects, e.g., task planning or causal relation discovery. Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI et al., 2023; Touvron et al., 2023b) have emerged with some reasoning capabilities  (Huang and Chang, 2023). However, there is observation that they still can not perform TR sufficiently well  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023), preventing their applications from solving complex real-world problems. In particular, TR requires a combination of various skills including mathematical and logical reasoning as well as commonsense knowledge  (Zhou et al., 2019; Qin et al., 2021).
@@
Figure 1: Our TG-LLM framework has two steps: text-to-temporal graph translation and temporal graph reasoning.
@@
Recent works mainly adopt general approaches to investigate and improve the TR capability of LLMs. For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark the leading LLMs on different TR tasks with standard Input/Output prompt, few-shot in-context learning (ICL) and Chain-of-Thought (CoT) reasoning  (Wei et al., 2023a). Similarly,  (Wei et al., 2023b) designs several specific types of prompts as prompt tuning.  (Li et al., 2023; Yuan et al., 2023) introduce pre-defined Python programs/rule-based templates to perform supervised fine-tuning (SFT). In addition,  (Tan et al., 2023a, b) adopt some extra strategies, which include specific pre-training, instruction tuning and reinforcement learning.
@@
Despite the effectiveness of such methods, they either ignore or not explicitly involve the intrinsic nature of TR. Humans perform complex TR on a timeline of events which are aligned with the entities and relations. These temporal concepts (e.g., ordering, duration, frequency, typical time) are then rigorously defined based on the timeline information. In other words, the aligned timeline (more generally, the temporal graph, TG) serves as a latent representation to help humans understand the patterns in TR. However, due to the lack of ground truth, the high-quality TG translation is a challenging task for most TR benchmarks. To solve this problem, we propose a synthetic dataset (TGQA), which is fully controllable and requires minimal supervision. We demonstrate the capability of TG translation learned on our dataset can be transferred to other TR tasks and benchmarks.
@@
Given a reliable TG, the key challenges of teaching LLMs TR include: (1) How can one introduce the necessary arithmetic and commonsense knowledge involved in TR? Prior works  (Lewis et al., 2021) show that explicitly introducing knowledge in context enhances the performance of LLMs. In this paper, we first identify all the valid time expressions, and then generate related knowledge (e.g., the temporal relation and time gap between the timestamps, the relative order of the gaps). (2) How can one teach LLM to perform deliberate reasoning? Generally, there exist two roadmaps: (i) translating natural language into logical statements, and using external symbolic engine for reasoning  Pan et al. (2023); (ii) using LLMs directly as the reasoning engine  (Zhu et al., 2023). For (i), the difficulty lies in accurate translation  Yang et al. (2023c) and the limited expressive power of formal logic. For (ii), there is no guarantee for the correctness of generated intermediate steps especially with insufficient training data  (Yang et al., 2023b). We adopt (ii) with the proposed bootstrapping method to generate reliable intermediate steps for supervised fine-tuning. We further improve the model performance with graph data augmentation which mitigates the data deficiency in TR tasks.
@@
  • We propose a new paradigm, TG-LLM, for language-based TR. In this framework, we first translate the context into a latent representation (temporal graph), and then perform reasoning on it. Extensive experiments prove that our novel approach results in superior performance compared to the baselines.
  • We design two approaches including Chain of Thought bootstrapping and graph data augmentation to teach LLM to generate consistent and faithful CoTs, which brings better performance than the vanilla CoT distillation.
  • We present a pipeline to create a synthetic dataset (TGQA) for question answering that requires TR. It is fully controllable and requires minimal supervision for text-temporal graph alignment. We show in experiments that fine-tuning on our dataset benefits LLM on other TR tasks and benchmarks.
@@
Relation <Event_B>? _and_Event_B_startTime>
@@
Facts When did the <Event_A> occur? <Event_A_startTime>
@@
Table 2: Reasoning types with the corresponding questions and answers in TGQA.
2 Dataset Construction
@@
In this section, we present the construction pipeline for TGQA dataset that is fully controllable and requires minimal supervision for text-temporal graph alignment. Compared with existing datasets  (Chen et al., 2021; Tan et al., 2023a), we have more diverse categories and types of TR questions (Table 2), and the ground-truth timeline. We first split the large temporal knowledge graph, YAGO11k (Dasgupta et al., 2018), into subgraphs with a restriction on the number of events, and anonymize the entities to avoid data leakage. Then each story is translated from the subgraph by GPT-3.5  (Ouyang et al., 2022). By using some rule-based templates, we obtain reliable question and answer (QA) pairs from the graph. Finally, to reduce the noise introduced from the misalignment between the subgraph and generated story, we propose a semi-automatic verification method.
@@
Step 1: Graph Splitting & Anonymization.  Existing temporal knowledge graphs (Leetaru and Schrodt, 2013; Dasgupta et al., 2018; García-Durán et al., 2018) usually have a large size. To facilitate the learning process, we split YAGO11k into subgraphs for story generation. Specifically, given a certain entity, we find its neighbors within three hops, and extract all the events happening between them. Since LLMs are supposed to do reasoning instead of memorization, we ensure that no overlapping exists between the events in training, validation and test sets. Additionally, we notice the data leakage problem of LLMs, i.e., prior knowledge of the test data has been implicitly obtained from the extensive pre-training. Thus, an anonymization strategy, i.e., changing entity names into random ones of the same type, is adopted. We generate a global mapping for entity names with GPT-3.5. To avoid confusion, we adopt obscure names that do not exist in YAGO11k.
@@
Step 2: Graph-based Open QA Creation.  In TGQA, each sample is in the form of (temporal graph, story, questions, answers) (Table 1). Based on the given subgraph, we generate a story and multiple QAs. We first ask GPT-3.5 to write a story based on the subgraph with the requirements to include all the events. It is observed that GPT-3.5 tends to ignore the end time of some events in the created story. To solve this problem, we separate the start and end time of the same event in prompt. On the other hand, to obtain a comprehensive benchmark, we consider all types of temporal reasoning, which include sequencing, duration, frequency, simultaneity, temporal relation, comparative analysis and facts extraction. Given these categories, we design multiple question types (Table 2) with a rule-based Python script to generate the corresponding Qs and As.
@@
Step 3: Quality Control.  In TGQA, we might introduce noise from the misalignment between the given subgraph and generated story. To reduce the noise, we propose a semi-automatic verification method. Fully manual inspection is expensive, but by first asking LLM we can narrow down potential errors, and only manually inspect the set of unanswered questions by LLM. Specifically, given a generated story, GPT-3.5 is queried on the time of each event in the graph. If it cannot give the correct answer, we consider the event as possibly missing in the story, which requires manual verification. We proved the effectiveness of our semi-automatic pipeline with fully manual verification of the test stories. Note that supervision is only required for story-TG alignment verification, since all the QAs are generated from rules. We show dataset statistics in Appendix A, and all the prompts involved in Appendix C.
@@ 3 TG-LLM
Motivated by human perception, we propose a new paradigm called TG-LLM. We first translate the text into a temporal graph (TG), and then guide the LLM to perform deliberate reasoning on it.
@@ 3.1 Text-to-TG Translation
Although LLM (with ICL) might have such capability, we observed a misalignment between the generated TG and pre-defined QAs, i.e, LLM making mistakes or focusing on irrelevant events. Since TG serves as the foundation of the following deliberate reasoning process, we provide a pipeline for high-quality TG data generation to fine-tune the LLM.
@@
Ground-truth TG Generation. For some datasets such as TGQA, a verified TG (corresponding to the story) is provided. We can directly fine-tune the LLM on this task with the ground truth. However, for most real-world applications, the biggest challenge is the lack of ground truth TG. We provide a pipeline for high-quality TG data generation. We first extract all the entities and relations from the QAs to help LLM focus on specific events. We then identify all the valid time expressions in the story. We provide these information to LLM for better alignment in TG construction, and verify the generated TG in a semi-automatic way.
(1) Entity & Relation Extraction: To obtain the entity and relation in pre-defined QAs, we consider two ways: parsing with rules or using GPT-3.5. Rule-based parsing, if applicable, is more efficient and reliable, which is prioritized in our experiments.
@@
(2) Temporal Info Identification: To identify all the valid time expressions, we first use GPT-3.5 to extract from the story, and then adopt a rule-based Python script for invalid output filtering and normalization. Besides bringing better alignment, the pre-processing of time expressions facilitates the introduction of external knowledge (their temporal relations and time gaps) into TR. We proved in experiments that explicitly introducing this information further improves model performance.
@@
(3) TG Construction & Verification: We generate the TG using GPT-3.5 with ICL. Specifically, we provide several in-context demonstrations which include a story and extracted entities, relations and time expressions as input and the corresponding TG as output (Table 15). For those stories without explicit time expressions, we skip Step (2), and ask GPT-3.5 to provide the temporal order of events. Similar to TGQA, we verify the generated TG by first asking GPT-3.5 the pre-defined QAs, and then manually checking if GPT-3.5 fails.
@@
Figure 2: In Chain of Thoughts (CoTs) bootstrapping, we accept CoTs leading to correct final answers and sample them according to their contrastive learning scores.
@@
Supervised Fine-tuning. We first conducted experiments to decide the best in-context format of TG. It is found out that providing a chronological list of events helps LLM perform better TR. In addition, it offers advantages to separate the start and end time of the same event. With these steps, the TG in context is transferred into a timeline with the alignment between entities, relations and times. We perform supervised fine-tuning (SFT) using Llama-2 model  (Touvron et al., 2023a) with Low-Rank Adaptation (LoRA)  (Hu et al., 2021). The input and output of the model are the story and aligned timeline, respectively. In experiments, we observe benefits from the SFT on our dataset to other temporal reasoning tasks.
3.2 Temporal Graph Reasoning
@@
Given the generated TGs, we teach LLM deliberate reasoning with SFT enhanced by CoT bootstrapping and graph data augmentation.
@@ 3.2.1 Bootstrapping Chain of Thoughts
It is observed that SFT on reliable CoTs brings better reasoning performance than that on standard Input/Output prompts  (Wang et al., 2023; Ho et al., 2023). Since asking humans to create the CoT data is not scalable, we use LLM to replace humans. This task is non-trivial for reasoning over knowledge graphs  (Saparov and He, 2023). In this section, we propose a bootstrapping pipeline, i.e., given a query, using LLM to generate several CoTs and selecting them as training data with a weighted sampling strategy. Compared with the conventional Best-of-N sampling, our proposal allows more training data diversity.
Specifically, we first prepare ICL examples for high-quality CoT generation. To facilitate learning, the pre-defined QAs are classified into different categories $\{Q_{1},Q_{2},\cdots$, $Q_{M}\}$. For each category $Q_{i}$, we randomly choose $N_{i}$ samples $\{(g_{j},e_{j},q_{j},a_{j})\}^{N_{i}}_{j=1}$, where $g$, $e$, $q$, $a$ denote TG, external knowledge, question and answer, respectively, and ask both GPT-3.5 and GPT-4 to provide diverse CoTs $\{c_{j}\}^{N_{i}}_{j=1}$. These CoTs will then be manually verified as ICL examples. Given a new training sample $(g_{j^{\prime}},e_{j^{\prime}},q_{j^{\prime}})$, we bootstrap $K$ CoTs with final answers $\{(c_{j^{\prime},k},\hat{a}_{j^{\prime},k})\}^{K}_{k=1}$ from GPT-3.5. We first refuse the CoTs leading to incorrect answers, i.e., $\hat{a}_{j^{\prime},k}\neq a^{*}_{j^{\prime}}$, where $a^{*}_{j^{\prime}}$ denotes the correct answer. For the accepted CoTs, we consider a weighted sampling strategy to balance usefulness and diversity. The sampling probability $P_{sample}(\cdot)$ is based on a contrastive learning $score$ (Eq. 1). The $score$ design, inspired by  (Wang et al., 2023), considers both perplexity $P(\cdot)$ of the correct answer and plausibility growth $G(\cdot)$ (Eq. 2). The definition of $G(\cdot)$ and $P(\cdot)$ are given in Eq. 3 and Eq. 4, respectively.
@@
where $c_{k}$ denotes the current CoT in consideration; $a^{\prime}$ denotes a certain wrong answer from the candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $t_{l}$ denotes the $l$-th token in $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.
3.2.2 Graph Data Augmentation
Compared with other tasks, reasoning suffers more from data insufficiency since more information (such as evidence, arguments, and logics) are involved in the intermediate steps  (Huang and Chang, 2023). To address this issue, we propose several graph data augmentation strategies (Figure 3).
Our framework involves two steps: text-to-TG translation and temporal knowledge graph reasoning. Notably, for the reasoning part, we train with ground-truth/verified TGs but infer on the estimated graphs. This discrepancy hurts the robustness of reasoning. Thus, we introduce some disturbances on the TGs during training. Note that the type of disturbances should be carefully designed in order to avoid confusing the LLM. We first investigate two types of disturbances: (i) remove irrelevant edges (Eq. 5) and (ii) replace edges by using relation synonyms (Eq. 6). An edge (event) is considered irrelevant if not involved in both QA and CoT.